OpenStudy (anonymous):

write a unit rate for the situation. driving 140 mi in 2h 45min (round to the nearest hundredth)

OpenStudy (anonymous):

\[140\div 2\tfrac{3}{4}\] \[140\div \frac{11}{4}\] \[140\times \frac{4}{11}\] \[\frac{560}{11}=50.91\] rounded